Two wires of different densities but same area of cross section are soldered together at one end and are stretched to a tension T. The velocity of a transverse wave in the first wire is double of that in the second wire. Find the ratio of the density of the first wire to that of second wire.

Grade:11

1 Answers

Deepak Patra
9 years ago
Sol. Let v1 = velocity in the 1st string ⇒ v1 = √(T /m1) Because m1 = mass per unit length = (1a1l1 / l1) = 1a1 where a1 = Area of cross section ⇒ v1 = √(T / p1a1) …(1) Let v2 = velocity in the second string ⇒ v2 = √(T /m2 ) ⇒ v2 = √(T / p2a2 ) …(2) Given that, v1 = 2v2 ⇒ √(T / p1a1) = 2 √(T / p2a2 ) ⇒ (T/a1p1) = 4(T/a2p2) ⇒ p1/p2 = 1/4 ⇒ p1 : p2 = 1 : 4 (because a1 = a2)